What is the force on an electron in an electric field of 500 N/C?

A charge of 5 X 10-6 C experiences a force of 10 N at a certain point in space. What is the electric field intensity?


1
Expert's answer
2022-03-08T09:40:51-0500

1.

"F=eE=80\cdot 10^{-18}~N."

2.

"E=\frac Fq=2~\frac{MN}C."